OpenStudy (anonymous):

Simplify (pi * R^3 + ( - pi * (R-r)^3)) /(Pi* R^2 + -pi * (R-r)^2)

OpenStudy (ash2326):

We have this \[\frac{\pi * R^3 + ( -\pi * (R-r)^3)} {\pi* R^2 + -\pi * (R-r)^2}\] let's take the \(\pi\) common from numerator and denominator and cancel it \[\frac{ R^3 - (R-r)^3} { R^2 - (R-r)^2}\] Do you get this?

OpenStudy (anonymous):

yea